Use GAUSS-JORDAN INVERSE METHOD to solve these systems of Linear Equations.

y-10z=-8

2х - бу=8

x+2z=7



1
Expert's answer
2021-06-18T12:14:32-0400

we are given

y-10z=-8

2х - бу=8

x+2z=7

this can be rewritten as

0x+y-10z=-8

2х - бу+0z=8

x+0y+2z=7

representing above equations in matrix form, we have

"\begin{pmatrix}
 0 & 1&-10 \\
 2& -6&0\\
1&0&2
\end{pmatrix}" "\begin{pmatrix}
 x \\
 y\\
z
\end{pmatrix}" ="\begin{pmatrix}
 -8 \\
 8\\
7
\end{pmatrix}"

let A="\begin{pmatrix}
 0 & 1&-10 \\
 2& -6&0\\
1&0&2
\end{pmatrix}" and b= "\begin{pmatrix}
 -8 \\
 8\\
7
\end{pmatrix}" and X="\begin{pmatrix}
x\\
 y\\
z
\end{pmatrix}"

find inverse of matrix A using GAUSS-JORDAN INVERSE METHOD by reducing the following matrix to row echelon form

= "\begin{pmatrix}
 0 & 1&-10 \mid&1&0&0\\
 2& -6& 0 \mid&0&1&0\\
1&0&2 \mid&0&0&1
\end{pmatrix}"

"R_1\leftrightarrow R_3"

= "\begin{pmatrix}
 1&0&2 \mid&0&0&1\\
 2& -6& 0 \mid&0&1&0\\
0 & 1&-10 \mid&1&0&0
\end{pmatrix}"

"R_2\to R_2-2R_1"

="\begin{pmatrix}
 1&0&2 \mid&0&0&1\\
 0& -6& -4\mid&0&1&-2\\
0 & 1&-10 \mid&1&0&0
\end{pmatrix}"

"R_2 \to- {1\over 6}R_2"

"R_3 \to 6R_3 +R_2"

="\begin{pmatrix}
 1&0&2 \mid&0&0&1\\
 0& 1& {4\over 6}\mid&0&-{1\over 6}&{2\over 6}\\
0 &0&-64 \mid&6&1&-2
\end{pmatrix}"

"R_3 \to -{1\over 64}R_3"

="\begin{pmatrix}
 1&0&2 \mid&0&0&1\\
 0& 1& {4\over 6}\mid&0&-{1\over 6}&{2\over 6}\\
0 &0&1 \mid&-{6\over 64}&-{1\over 64}&{2\over 64}
\end{pmatrix}"

"R_2 \to R_2-{4\over 6}R_3"

"R_1 \to R_1-2R_3"


="\begin{pmatrix}
 1&0&0 \mid&{12\over 64}&{2\over 64}&{15\over 16}\\
 0& 1& 0\mid&{4\over 64}&-{5\over 32}&{5\over1 6}\\
0 &0&1 \mid&-{6\over 64}&-{1\over 64}&{2\over 64}
\end{pmatrix}"


"\implies" "A^{-1}" = "\begin{pmatrix}
 {12\over 64}&{2\over 64}&{15\over 16}\\
 {4\over 64}&-{5\over 32}&{5\over 16}\\
-{6\over 64}&-{1\over 64}&{2\over 64}
\end{pmatrix}"


A-1AX=A-1b


"\begin{pmatrix}
 {12\over 64}&{2\over 64}&{15\over 16}\\
 {4\over 64}&-{5\over 32}&{5\over 16}\\
-{6\over 64}&-{1\over 64}&{2\over 64}
\end{pmatrix}" "\begin{pmatrix}
 0 & 1&-10 \\
 2& -6&0\\
1&0&2
\end{pmatrix}" "\begin{pmatrix}
x\\
 y\\
z
\end{pmatrix}" ="\begin{pmatrix}
 {12\over 64}&{2\over 64}&{15\over 16}\\
 {4\over 64}&-{5\over 32}&{5\over 16}\\
-{6\over 64}&-{1\over 64}&{2\over 64}
\end{pmatrix}" "\begin{pmatrix}
 -8 \\
 8\\
7
\end{pmatrix}"


"\begin{pmatrix}
 1& 0&0 \\
 0& 1&0\\
0&0&1
\end{pmatrix}" "\begin{pmatrix}
x\\
 y\\
z
\end{pmatrix}" ="\begin{pmatrix}
5.31250\\
 0.43750\\
0.84375
\end{pmatrix}"



"\therefore" "x=5.31250"

"y=0.43750"

"z=0.84375"